fontimport(1) teTeX fontimport(1) NAME fontimport−importfontsintotheteTeXdirectorystructure SYNOPSIS fontimport[−d][−f][−s][−t][−mMF−mode]path... DESCRIPTION fontimportimportsTFMandPKfontfilesintoappropriateplacesintheteTeXdirectorytreeand revisesteTeX’sfilenamedatabaseaccordingly. Eachpathmaybeafontfileoradirectory,eitherabsoluteorrelativetothecurrentdirectory. Allsubdirectoriesofagivendirectoryaresearchedforfontfiles.Don’tspecifydirectorynames thataresymboliclinks,however,since(atleastsomeversionsof)find(1)won’tdereferencethem. Afont’sdestinationisdeterminedbyvariablesinthefiletexmf.cnf.Bydefault,fontimportwill issueawarningtostderrinsteadofoverwritinganexistingfontfile.The−fflagforcesoverwriting. fontimportcancopewithfontnamesconformingthe8.3−namescheme,too,soyoucanimport fontsfromeitheraTDS-compliantorMSDOSinstallation.emTeX’sfontlibrariesarenotsup-ported.Ifyouwanttoimportfontsfromafontlibrary,youneedtounpackitwithemTeX’s fontlibutilityfirst. Thefullpathnameofeachfontsuccessfullyimportedisechoedtostdout.Whenfontimportcan notparseafont’sname,itwillimportthefontusingapathelementof"/unknown/"andissuea warningonstderr. fontimporttriestoguessthemodeofPKfiles.Withthe-mflagyoumayspecifyafallbackmode tobeusedifthemodecan’tbeguessed. Options: −mMF-MODE defaultMetafontmodeforPKfilesifautodetectionfails −ddeletefontfilesaftercopying(i.e.,movethem) −fforceoverwriteexistingfonts −sdonotuseotherdestinationsthanthesystem’smaintexmftree −ttestmode:justechoeachfont’sdestination(ifany);don’tmoveorcopy SEEALSO allec(1),allneeded(1),fontexport(1),Fontname:FilenamesforTeXFonts (http://tug.org/fontname). AUTHOR ThomasEsser(te@informatik.uni−hannover.de) teTeX 4/99 1